# Semantic MIMO Systems for Speech-to-Text Transmission

### v1

@@ Abstract
Semantic communications have been utilized to execute numerous intelligent tasks by transmitting task-related semantic information instead of bits. In this article, we propose a semantic-aware speech-to-text transmission system for the single-user multiple-input multiple-output (MIMO) and multi-user MIMO communication scenarios, named SAC-ST. Particularly, a semantic communication system to serve the speech-to-text task at the receiver is first designed, which compresses the semantic information and generates the low-dimensional semantic features by leveraging the transformer module. In addition, a novel semantic-aware network is proposed to facilitate the transmission with high semantic fidelity to identify the critical semantic information and guarantee it is recovered accurately. Furthermore, we extend the SAC-ST with a neural network-enabled channel estimation network to mitigate the dependence on accurate channel state information and validate the feasibility of SAC-ST in practical communication environments. Simulation results will show that the proposed SAC-ST outperforms the communication framework without the semantic-aware network for speech-to-text transmission over the MIMO channels in terms of the speech-to-text metrics, especially in the low signal-to-noise regime. Moreover, the SAC-ST with the developed channel estimation network is comparable to the SAC-ST with perfect channel state information.
@@ I Introduction
Semantic communications are the second level of communications defined by Shannon and Weaver [2] to deliver the semantic exchange and optimize the communication systems at the semantic level. To alleviate the scarcity of bandwidth resources and the explosion of data traffic, the semantic communication problem has been listed as one of the promising techniques for future intelligent communications [3, 4, 5].
@@
The conventional communications, i.e., the first level of communications, convert the input message into a bit sequence and aim to achieve a low bit-error rate (BER) by leveraging proper coding and decoding algorithms. The bit-oriented transmission framework has undergone a thriving development, but the attention to semantic information has been neglected. The ultimate goal of semantic communications is to exploit the semantic information inside the source and facilitate the semantic transmission to convey the meaning of the transmitter to the receiver. However, due to the limitation of a unified theory to represent the semantics of different sources and a mathematical formula to quantify the semantic information, the investigation of semantic communications has experienced a few decades of stagnation. Thanks to the success of artificial intelligence (AI) in recent years, a deep learning (DL)-enabled semantic communication architecture has been proposed and has shown its great potential to tackle the technical challenges across many aspects of conventional communications, e.g., resource allocation [6] and physical layer communications [7]. Such success is attributed to the learning and fitting capabilities of diverse neural networks, which break the constraints of a mathematical model to exchange semantic information.
@@
The DL-enabled semantic communication paradigm requires the least semantic ambiguity between the source and the recovered messages. It maps the source message into the low-dimensional semantic features, which reduces the volume of transmission data and mitigates the network traffic without performance degradation. Inspired by this, a preliminary research work of DL-enabled semantic communications for text transmission has been developed in [8]. This work designed a transformer-powered joint semantic-channel codecs mechanism, named DeepSC, to minimize the semantic error by measuring the meaning difference between the input and recovered sentences. Inspired by DeepSC, Peng et al. [9] further proposed R-DeepSC to improve the semantic fidelity and against the semantic noise inherent in the sentence through a semantic error corrector. In [10], a text semantic communication system is developed to ensure transmission reliability by adopting the hybrid automatic repeat request (HARQ) scheme. Liang et al. [11] studied a reasoning-based semantic communication paradigm to represent semantics through a graph-based knowledge architecture and a semantic interpretation framework. Xie et al. investigated task-oriented semantic communications for text transmission in [12] and introduced a memory component in [13] to tackle the transmission problems in dynamic channel environments. Nam et al. [14] developed a novel sequential semantic communication system for text-to-image generation through multi-modal generation techniques. Moreover, a DL-enabled semantic communication system for speech signals, named DeepSC-S, has been designed in [15] by extracting and transmitting global semantic information. Han et al. [16] proposed an efficient speech semantic communication system by introducing the redundancy removal module to drop insignificant semantics and reduce the transmission data. In [17], a deep speech semantic transmission system is developed to further improve the intelligibility of the reconstructed speech by leading to a flexible rate-distortion trade-off. Weng et al. [18] proposed a task-oriented semantic communication system for speech transmission, named DeepSC-ST, which dramatically outperforms the conventional speech transmission paradigm. In [19], a unified semantic communication system for multimodal data has been proposed by fusing shared trainable parameters amongst various tasks to control the dimension of extracted semantics and reduce the transmission data.
@@
At the same time, semantic communications for image/video transmission have attracted numerous research interests[20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Particularly, a generative adversarial network (GAN)-enabled semantic communication system for image transmission has been proposed in [20] by interpreting the image meaning. Xu et al. [21] developed a joint image transmission and scene classification scheme in a reinforcement learning-enabled semantic communication system. A robust image semantic transmission scheme has been studied in [22] by incorporating a masked vector quantized-variational autoencoder to combat semantic error and improve image classification accuracy. Nan et al. [23] designed a training efficiency mechanism for image semantic communication by utilizing a distributed learning architecture to balance the trade-off between computation capability and user privacy. A video semantic transmission network is established in [26], which utilizes a semantic detector at the receiver to reduce the semantic error. Thanks to the advancements in computer vision, the semantic communication for performing video tasks, extended reality (XR), has been designed in [29], in which the insignificant semantic information is highly compressed to mitigate the transmission traffic.
@@
DL-enabled semantic communications have also been investigated in multiple-input multiple-output (MIMO) transmission scenarios[30, 31, 32, 33, 34, 35]. Wu et al. [30] proposed a vision transformer-enabled semantic communication system for image transmission over MIMO channels by using a self-attention module to exploit the contextual semantic features and the channel conditions. In [31], an orthogonal space-time block code (OSTBC) is utilized to improve the image transmission robustness to MIMO channel variants, and an equalizer is leveraged to estimate the codewords. Yao et al. [32] designed a versatile semantic coded transmission scheme over MIMO fading channels to adaptively adjust the coding rate allocation and stream mapping. Additionally, Luo et al. [33] considered a semantic precoding-aided MIMO system to alleviate the effect of physical channels by a new multimodal information fusion scheme. A massive MIMO semantic transmission system for image restoration is developed in [34], in which the post-filter and image statistics are leveraged to improve the image quality. In [35], Zhang et al. established a channel feedback-enabled semantic communication framework in MIMO systems to ensure the image reconstruction and proposed a metric, termed semantic distortion outage probability, to measure the reliability of semantic transmission.
@@
In this article, we design a semantic-aware speech-to-text transmission scheme in single-user MIMO (SU-MIMO) and multi-user MIMO (MU-MIMO) communication systems, named SAC-ST. We argue that previous research works on semantic communications for speech have limited contributions to the analysis of semantic information, and the attention mechanisms for extracting the semantic features are not fully exploited. To address these challenges, a transformer-enabled semantic communication system for speech-to-text transmission over MIMO channels, named DeepSC-ST 2.0, is developed, which removes the semantic redundancy and improves the system performance compared to the DeepSC-ST [18]. Moreover, a semantic-aware network is proposed to boost the transmission with high semantic fidelity by identifying and recovering the critical semantic information. The contributions of this paper are summarized as follows:
@@
  • A semantic communication system for speech-to-text transmission over MIMO fading channels, named DeepSC-ST 2.0, is developed by utilizing the transformer to squeeze the input speech and extract the low-dimensional semantic features. The proposed scheme achieves superior speech-to-text performance and serves users only requesting text information.
@@
  • To improve the semantic fidelity, we propose a novel neural network-enabled semantic-aware network incorporated with the DeepSC-ST 2.0 to form the proposed SAC-ST. The semantic-aware network perceives the importance of different semantic information and grants high priority to the essential semantic information to be recovered accurately.
  • The SAC-ST is adaptively deployed in the SU-MIMO and MU-MIMO communication systems. To validate the feasibility of SAC-ST in practical communication scenarios, a MIMO channel estimation network, named ChanEst network, is established and integrated into the SAC-ST as a proof-of-concept in the SU-MIMO and MU-MIMO transmission paradigms.
@@
  • Simulation results verify the superiority of the SAC-ST compared to the framework without the semantic-aware network. Moreover, the SAC-ST with the ChanEst network achieves remarkable performance similar to the SAC-ST with the perfect channel state information.
@@ II Motivations
This section briefly describes the MIMO transmission system with singular value decomposition (SVD) precoding. We will then discuss the motivations of the proposed semantic-aware communication system over SU-MIMO channels and MU-MIMO channels.
@@ II-A MIMO Transmission with SVD Precoding
The SVD precoding-based MIMO channel can be treated as multiple independent SISO channels, and the noise SNRs of these SISO channels are inconsistent due to the difference of singular values, denoted as
@@
where the first SISO channel has the highest SNR value of $\left(\Lambda_{11}\right)^{2}\ast\sigma$ and exhibits minimal noise impairment.
@@ II-B Motivations of MIMO Semantic Transmission
In conventional communication systems, the input message is mapped into a bit sequence, and the bits transmitted over different SISO channels are considered equally important. However, the semantic communication paradigm converts the input message into semantic features. The semantic information contained by different semantic features may vary, which indicates the importance of different semantic features to the intelligent tasks is unequally distributed. Inspired by this, we select the semantic features with high importance and place them into the SISO channels with high SNRs. By doing so, the essential semantic features are recovered accurately at the receiver, which maximizes the semantic fidelity and reduces errors in the recovered message. An example of the SU-MIMO semantic communication system with SVD precoding is shown in Fig. 2. From the figure, the key information “Leo” is placed into the top SISO channel with the highest SNR, while the less critical information “am” is transmitted over the SISO with the lowest SNR. By doing so, the sentence restored by the semantic transmission paradigm over MIMO channels with SVD precoding is more intelligible.
@@
To achieve the transmission mechanism that prioritizes semantic importance, the precoding operation performed to symbols $\boldsymbol{x}$ is responsible for eliminating the MUI and enabling the MIMO channel decomposition to attain equivalent multiple SISO channels. The conventional block diagonalization (BP) algorithm [36] is proficient in accomplishing this objective. However, it demands a considerable computational burden due to the necessity of two SVD precoding operations and requires each receiver user to possess knowledge of the channel state information (CSI) pertaining to all other users. Therefore, we propose a novel precoding algorithm to overcome the limitations of the BP algorithm. Particularly, the MIMO channel ${\boldsymbol{H}}_{k}$ is decomposed according to the SVD operation as follows,
@@
By doing so, the MUI is completely removed and the MIMO channel, ${\boldsymbol{H}}_{k}$, behaves as multiple independent SISO channels to deliver the information of the $k$-th user, which inspires to assign the essential information to the good SISO channels to achieve the high semantic fidelity transmission. However, when transmitting the symbols of the $k$-th user through (11), the precoded symbols of other users still undergo the channel fading and the MUI. To ensure overall semantic fidelity for $K$ users, we transmit the insignificant information of other users when the objective is to recover the information of the $k$-th user accurately.
@@ III Semantic-Aware Speech-to-Text Transmission Over MIMO Systems with Channel Estimation
In this section, the details of the proposed SAC-ST are introduced. We consider the SU-MIMO and the MU-MIMO transmission schemes. Besides, a channel estimation network is developed to eliminate the reliance on accurate CSI and deployed in both the SU-MIMO and MU-MIMO communication scenarios.
@@ III-A SU-MIMO Communicaton System
During the SU-MIMO transmission, the data transmitted over multiple SISO channels is aggregated to one user at the receiver. In the proposed SAC-ST, we design the DeepSC-ST 2.0 by utilizing the transformer module to serve the speech-to-text task and develop a semantic-aware network to quantify the importance of the semantic features extracted by the semantic encoder, as shown in Fig. 3. From the figure, the first and second training stages are performed to train the DeepSC-ST 2.0 and the semantic-aware network, respectively, which are adopted in the testing stage to achieve the semantic transmission with high semantic fidelity by identifying the essential semantic features and transmitting them over the SISO channels with high SNRs.

### v2

@@ Abstract
Semantic communications have been utilized to execute numerous intelligent tasks by transmitting task-related semantic information instead of bits. In this article, we propose a semantic-aware speech-to-text transmission system for the single-user multiple-input multiple-output (MIMO) and multi-user MIMO communication scenarios, named SAC-ST. Particularly, a semantic communication system to serve the speech-to-text task at the receiver is first designed, which compresses the semantic information and generates the low-dimensional semantic features by leveraging the transformer module. In addition, a novel semantic-aware network is proposed to facilitate transmission with high semantic fidelity by identifying the critical semantic information and guaranteeing its accurate recovery. Furthermore, we extend the SAC-ST with a neural network-enabled channel estimation network to mitigate the dependence on accurate channel state information and validate the feasibility of SAC-ST in practical communication environments. Simulation results will show that the proposed SAC-ST outperforms the communication framework without the semantic-aware network for speech-to-text transmission over the MIMO channels in terms of the speech-to-text metrics, especially in the low signal-to-noise regime. Moreover, the SAC-ST with the developed channel estimation network is comparable to the SAC-ST with perfect channel state information.
@@ I Introduction
Semantic communications are the second level of communications defined by Shannon and Weaver [2] to deliver the semantic exchange and optimize the communication systems at the semantic level. To alleviate the scarcity of bandwidth resources and the explosion of data traffic, the semantic communication problem has been listed as one of the promising techniques for future intelligent communications [3, 4, 5, 6].
@@
The conventional communications, i.e., the first level of communications, convert the input message into a bit sequence and aim to achieve a low bit-error rate (BER) by leveraging proper coding and decoding algorithms. The bit-oriented transmission framework has undergone a thriving development, but the attention to semantic information has been neglected. The ultimate goal of semantic communications is to exploit the semantic information inside the source and facilitate the semantic transmission to convey the meaning of the transmitter to the receiver. However, due to the limitation of a unified theory to represent the semantics of different sources and a mathematical formula to quantify the semantic information, the investigation of semantic communications has experienced a few decades of stagnation. Thanks to the success of artificial intelligence (AI) in recent years, the deep learning (DL)-enabled semantic communication architecture has been proposed and has shown its great potential to tackle the technical challenges across many aspects of conventional communications, e.g., resource allocation [7] and physical layer communications [8]. Such success is attributed to the learning and fitting capabilities of diverse neural networks, which break the constraints of a mathematical model to exchange semantic information.
@@
The DL-enabled semantic communication paradigm requires the least semantic ambiguity between the source and the recovered messages. It maps the source message into the low-dimensional semantic features, which reduces the volume of transmission data and mitigates the network traffic without performance degradation. Inspired by this, a preliminary research work of DL-enabled semantic communications for text transmission has been developed in [9]. This work designed a transformer-powered joint semantic-channel codecs mechanism, named DeepSC, to minimize the semantic error by measuring the meaning difference between the input and recovered sentences. Inspired by DeepSC, Peng et al. [10] further proposed R-DeepSC to improve the semantic fidelity and against the semantic impairments inherent in the sentence through a semantic error corrector. In [11], a text semantic communication system is developed to ensure transmission reliability by adopting the hybrid automatic repeat request (HARQ) scheme. Liang et al. [12] studied a reasoning-based semantic communication paradigm to represent semantics through a graph-based knowledge architecture and a semantic interpretation framework. Xie et al. investigated task-oriented semantic communications for text transmission in [13] and introduced a memory component in [14] to tackle the transmission problems in dynamic channel environments. Nam et al. [15] developed a novel sequential semantic communication system for text-to-image generation through multi-modal generation techniques. Moreover, a DL-enabled semantic communication system for speech signals, named DeepSC-S, has been designed in [16] by extracting and transmitting global semantic information. Wei et al. [17] devised a real-time audio semantic communication system to strengthen the audio quality and alleviate the physical channel attenuation by capturing the long-distance dependency. Han et al. [18] proposed an efficient speech semantic communication system by introducing the redundancy removal module to drop insignificant semantics and reduce the transmission data. In [19], a deep speech semantic transmission system is developed to further improve the intelligibility of the reconstructed speech by leading to a flexible rate-distortion trade-off. Weng et al. [20] proposed a task-oriented semantic communication system for speech transmission, named DeepSC-ST, which dramatically outperforms the conventional speech transmission paradigm. In [21], a unified semantic communication system for multimodal data has been proposed by fusing shared trainable parameters amongst various tasks to control the dimension of extracted semantics and reduce the transmission data.
@@
At the same time, semantic communications for image/video transmission have attracted numerous research interests[22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34]. Particularly, a generative adversarial network (GAN)-enabled semantic communication system for image transmission has been proposed in [22] by interpreting the image meaning. Xu et al. [23] developed a joint image transmission and scene classification scheme in a reinforcement learning-enabled semantic communication system. A robust image semantic transmission scheme has been studied in [24] by incorporating a masked vector quantized-variational autoencoder to combat semantic error and improve image classification accuracy. Nan et al. [25] designed a training efficiency mechanism for image semantic communication by employing a distributed learning architecture to balance the trade-off between computation capability and user privacy. More recently, Cicchetti et al. [28] established a semantic communication framework that combines text with compressed image embeddings and leverages a diffusion model to reconstruct images. Zhao et al. [29] developed a large language model-powered semantic communication system to improve the reconstructed image quality by transmitting multimodal features and Qiao et al. [30] introduced a latency-aware semantic communication mechanism to enable low-rate and low-latency visual transmission with high semantic fidelity. A video semantic transmission network is established in [31], which utilizes a semantic detector at the receiver to reduce the semantic error. Thanks to the advancements in computer vision, the semantic communication for performing video tasks, extended reality (XR), has been designed in [34], in which the insignificant semantic information is highly compressed to mitigate the transmission traffic.
@@
DL-enabled semantic communications have also been investigated in multiple-input multiple-output (MIMO) transmission scenarios[35, 36, 37, 38, 39, 40]. Wu et al. [35] proposed a vision transformer-enabled semantic communication system for image transmission over MIMO channels by using a self-attention module to exploit the contextual semantic features and the channel conditions. In [36], an orthogonal space-time block code (OSTBC) is utilized to improve the image transmission robustness to MIMO channel variants, and an equalizer is leveraged to estimate the codewords. Yao et al. [37] designed a versatile semantic coded transmission scheme over MIMO fading channels to adaptively adjust the coding rate allocation and stream mapping. Additionally, Luo et al. [38] considered a semantic precoding-aided MIMO system to alleviate the effect of physical channels by a new multimodal information fusion scheme. A massive MIMO semantic transmission system for image restoration is developed in [39], in which the post-filter and image statistics are leveraged to improve the quality of reconstructed images. In [40], Zhang et al. established a channel feedback-enabled semantic communication framework in MIMO systems to ensure the image reconstruction and proposed a metric, termed semantic distortion outage probability, to measure the reliability of semantic transmission.
@@
In this article, we devise a semantic-aware speech-to-text transmission scheme in single-user MIMO (SU-MIMO) and multi-user MIMO (MU-MIMO) communication systems, named SAC-ST. We argue that previous state-of-the-art (SOTA) works on semantic communications for speech transmission, e.g., SemAudio [17] and DeepSC-ST [20], have limited contributions to the analysis of semantic information, and the attention mechanisms for extracting the semantic features are not fully exploited. To address these challenges, a transformer-enabled semantic communication system for speech-to-text transmission over MIMO channels, named DeepSC-ST 2.0, is first developed. The DeepSC-ST 2.0 is evolved from the DeepSC-ST. Particularly, in DeepSC-ST 2.0, we redesign the semantic coding mechanism to avoid the intermediate spectrum and replace recurrent neural network (RNN) modules in DeepSC-ST with transformer modules, further eliminating the semantic redundancy and exploring the long-dependence of semantic information to improve the system performance compared to the DeepSC-ST, especially when contending with lengthy speech sequences. Moreover, based on the singular value decomposition (SVD) precoding of MIMO channels, we investigate the inequality of semantic information and utilize the SNR difference of decomposed equivalent SISO subchannels to boost the high semantic fidelity transmission by transmitting critical semantics over SISO subchannels with high SNRs. A novel semantic-aware network tailored for such a semantic transmission mechanism is proposed to identify the importance of different semantics and return the corresponding coefficients. The contributions of this paper are summarized as follows:
@@
  • A semantic communication system for speech-to-text transmission over MIMO fading channels, named DeepSC-ST 2.0, is developed to serve users only requesting text information by leveraging the convolutional neural network (CNN) to condense the input speech sequence and the transformer to extract the low-dimensional semantic features. The proposed scheme avoids the additional spectrum extraction and achieves superior speech-to-text transmission performance compared to the existing literature [20, 18].
@@
  • To improve the semantic fidelity, we propose a novel neural network-enabled semantic-aware network, which is incorporated with DeepSC-ST 2.0 to form the proposed SAC-ST. According to the inequality of semantic features and the inconsistent SNRs of multiple independent SISO subchannels based on the SVD precoding of MIMO channels, the semantic-aware network perceives the essential semantics and grants them high priority to be recovered accurately over good subchannels.
  • To validate the feasibility of SAC-ST in practical communication scenarios, a novel channel precoding algorithm is carried out to reduce the computational complexity in the MU-MIMO environment, and a lightweight MIMO channel estimation network is jointly designed with the SAC-ST to minimize the dependence on perfect channel state information (CSI) in the SU-MIMO and MU-MIMO transmission paradigms.
@@
  • Simulation results verify the superiority of the SAC-ST compared to the framework without the semantic-aware network. Moreover, the SAC-ST with the ChanEst network achieves remarkable performance similar to the SAC-ST with the perfect channel estimation.
@@ II Motivations
This section briefly describes the MIMO transmission system with SVD precoding. We will then discuss the motivations of the proposed semantic-aware communication system over SU-MIMO channels and MU-MIMO channels.
@@ II-A MIMO Transmission with SVD Precoding
The SVD precoding-based MIMO channel can be treated as multiple independent SISO subchannels, and the noise SNRs of these SISO subchannels are inconsistent due to the difference of singular values, denoted as
@@
where the first SISO subchannel has the highest SNR value of $\left(\Lambda_{11}\right)^{2}\ast\sigma$ and exhibits minimal noise impairment.
@@ II-B Motivations of MIMO Semantic Transmission
In conventional communication systems, the input message is mapped into the bit sequence, as shown in Fig. 2 (a). From the figure, the bits transmitted over different SISO subchannels are considered equally important and contain no semantics. In this context, the key information “Leo” may be corrupted in the recovered message. The semantic communication paradigm converts the input message into semantic features and the information contained by different semantic features is inconsistent, indicating the importance of various features is unequally distributed. Inspired by this, the semantic features with high importance are identified and placed into the SISO subchannels with high SNRs prior to transmission. By doing so, the essential semantic information is received accurately to maximize the semantic fidelity of the recovered message. As shown in Fig. 2 (b), the semantic features corresponding to the key information “Leo” are placed into the top SISO subchannel with the highest SNR, while the semantic features corresponding to the less critical information “am” are transmitted over the SISO subchannel with the lowest SNR, which preserves the key information “Leo” and ensures the intelligibility of the obtained sentence. However, an efficient algorithm to distinguish the importance of the extracted semantic features and a sort operation to these semantic features are necessary to perform such a semantic transmission mechanism.
@@
To achieve the transmission mechanism that prioritizes semantic importance, the precoding operation performed to symbols $\boldsymbol{x}$ is responsible for eliminating the MUI and enabling the MIMO channel decomposition to attain equivalent multiple SISO subchannels. The conventional block diagonalization (BP) algorithm [41] is proficient in accomplishing this objective. However, it requires each receiver user to possess knowledge of the CSI pertaining to all other users and demands a considerable computational burden due to the necessity of two SVD precoding operations, including the first one to obtain a precoding matrix, $\boldsymbol{V}^{\mathrm{B}}$, that completely removes the MUI and the second one to decompose the channel of each user into parallel subchannels according to a precoding matrix $\boldsymbol{V}^{\mathrm{P}}$. Then the joint precoding matrix of the BP algorithm is $\boldsymbol{V}^{\mathrm{BP}}=\boldsymbol{V}^{\mathrm{B}}\boldsymbol{V}^{\mathrm{P}}$. Toward this end, we propose a novel precoding algorithm to overcome the limitations of the BP algorithm. Particularly, the MIMO channel ${\boldsymbol{H}}_{k}$ is decomposed according to the SVD operation as follows,
@@
By doing so, the MUI is completely removed and the MIMO channel, ${\boldsymbol{H}}_{k}$, behaves as multiple independent SISO subchannels to deliver the information of the $k$-th user, which inspires to assign the essential information to the good SISO subchannels to achieve the high semantic fidelity transmission. However, when transmitting the symbols of the $k$-th user through (11), the precoded symbols of other users still undergo the channel fading and the MUI. To ensure overall semantic fidelity for $K$ users, we transmit the insignificant information of other users when the objective is to recover the information of the $k$-th user accurately.
@@ III Semantic-Aware Speech-to-Text Transmission Over MIMO Systems with Channel Estimation
This section introduces the details of the proposed SAC-ST. We consider the SU-MIMO and MU-MIMO transmission schemes. Besides, a channel estimation network is developed to eliminate the reliance on accurate CSI and deployed in both the SU-MIMO and MU-MIMO communication scenarios.
@@ III-A SU-MIMO Communicaton System
During the SU-MIMO transmission, the data transmitted over multiple SISO subchannels is aggregated to one user at the receiver. In the proposed SAC-ST, we design the DeepSC-ST 2.0 to serve the speech-to-text task. It is noteworthy that the existing works [20, 18] require a pre-processing step to input speech signals before neural networks, which runs counter to intelligent semantic extraction and hinders the end-to-end gradient backpropagation to the original speech signals. To address this, the one-dimensional (1D) CNN is adopted to directly condense speech signals into intermediate representations, and the transformer is leveraged to explore the semantics inherent in these representations, as shown in Fig. 3. Additionally, a semantic-aware network is developed to quantify the importance of the semantic features extracted by the semantic encoder, as shown in Fig. 4. From the figure, the first and second training stages are performed to train the DeepSC-ST 2.0 and the semantic-aware network, respectively, which are adopted in the testing stage to achieve the semantic transmission with high semantic fidelity by identifying the essential semantic features and transmitting them over the SISO subchannels with high SNRs.
III-A1 First training stage
The goal of the first training stage is to train the DeepSC-ST 2.0 that recovers the text sequence at the receiver from the input speech over MIMO channels with SVD precoding. As shown in Fig. 4 (a), the 1D CNN and transformer-enabled semantic encoder converts the input speech, $\boldsymbol{s}$, into the low-dimensional semantic features, $\boldsymbol{f}$. The channel encoder processes $\boldsymbol{f}$ and produces the symbols, $\boldsymbol{x}$, that are precoded into $\widetilde{\boldsymbol{x}}$ to be transmitted over the MIMO channel, $\boldsymbol{H}$. Therefore, $\widetilde{\boldsymbol{x}}$ can be denoted as
Figure 3: The model structure of (a) the semantic encoder in the existing works (b) and the proposed semantic encoder.